# Generating, Fast and Slow: Scalable Parallel Video Generation with Video Interface Networks

### v1

@@ Abstract
Diffusion Transformers (DiTs) can generate short photorealistic videos, yet directly training and sampling longer videos with full attention across the video remains computationally challenging. Alternative methods break long videos down into sequential generation of short video segments, requiring multiple sampling chain iterations and specialized consistency modules. To overcome these challenges, we introduce a new paradigm called Video Interface Networks (VINs), which augment DiTs with an abstraction module to enable parallel inference of video chunks. At each diffusion step, VINs encode global semantics from the noisy input of local chunks and the encoded representations, in turn, guide DiTs in denoising chunks in parallel. The coupling of VIN and DiT is learned end-to-end on the denoising objective. Further, the VIN architecture maintains fixed-size encoding tokens that encode the input via a single cross-attention step. Disentangling the encoding tokens from the input thus enables VIN to scale to long videos and learn essential semantics. Experiments on VBench demonstrate that VINs surpass existing chunk-based methods in preserving background consistency and subject coherence. We then show via an optical flow analysis that our approach attains state-of-the-art motion smoothness while using 25-40% fewer FLOPs than full generation. Finally, human raters favorably assessed the overall video quality and temporal consistency of our method in a user study. All videos shown in this manuscript can be found here.
@@ 1 Introduction
Diffusion Transformers (DiTs) [34] have heralded a revolution in generative modeling and have been scaled up to generate images [5], videos [3, 9, 44, 54], and 3D assets [31] at unprecedented fidelity. Training DiTs on longer input contexts yields remarkable generalization, albeit imposing quadratic computational and memory bottlenecks. The implications are particularly apparent in the case of high-dimensional modalities like video, where training and inference with DiTs on long videos ($>$ 64 frames) are slow and expensive [9] (see Fig 1(a), Full Denoising). Furthermore, base models trained on short videos often show motion stagnation and repetition when extended to synthesize longer videos [2, 17]. Practitioners circumvent these constraints by generating video frames in temporal chunks using the previously generated frames as a context window for the next chunk [19, 1] proceeding in a left-to-right sequence (Fig. 1(a), Autoregressive Denoising)
@@ 3.3 Learning Objective
where $X_{t}^{i}\equiv X_{t}^{iN_{s}:(i+1)N_{s}}$ and $\epsilon_{t}^{i}\equiv\epsilon_{t}^{iN_{s}:(i+1)N_{s}}$ denote the $i^{th}$ chunk of the input and noise estimate, respectively. Eq. (8) enables the DiT to simultaneously train on different temporal chunks in parallel. Moreover, end-to-end training enables the architectures to bootstrap off each other where (1) the DiT learns to ground denoising in encoded tokens and (2) the VIN learns to refine global tokens to enable downstream denoising. In practice, neighboring chunks have useful local contexts for denoising. During training, for each chunk $i$, we also append $N_{local}$ tokens from the last $F_{local}$ frames of the previous chunk $i-1$ to the DiT.
@@ 4.6 User Study
We set up a user study (design outlined in Appendix 12) where a cohort of humans was asked to compare videos generated by VINs against other methods on (1) overall appearance and (2) temporal consistency. Fig. 12 reports the percentage results with which VIN is deemed better, comparable, or worse. In contrast to the imaging quality results of Section 4.3, humans rated VIN-generated videos to be comparable on average or even better (loss rate $<$ 30 %). Moreover, the judged temporal consistency largely correlates with the VBench results, where VIN preference matches full generation, substantially improves upon parallel methods, and outperforms sequential generation.

### v2

@@ Abstract
Diffusion Transformers (DiTs) can generate short photorealistic videos, yet directly training and sampling longer videos with full attention across the video remains computationally challenging. Alternative methods break long videos down into sequential generation of short video segments, requiring multiple sampling chain iterations and specialized consistency modules. To overcome these challenges, we introduce a new paradigm called Video Interface Networks (VINs), which augment DiTs with an abstraction module to enable parallel inference of video chunks. At each diffusion step, VINs encode global semantics from the noisy input of local chunks and the encoded representations, in turn, guide DiTs in denoising chunks in parallel. The coupling of VIN and DiT is learned end-to-end on the denoising objective. Further, the VIN architecture maintains fixed-size encoding tokens that encode the input via a single cross-attention step. Disentangling the encoding tokens from the input thus enables VIN to scale to long videos and learn essential semantics. Experiments on VBench demonstrate that VINs surpass existing chunk-based methods in preserving background consistency and subject coherence. We then show via an optical flow analysis that our approach attains state-of-the-art motion smoothness while using 25-40% fewer FLOPs than full generation. Finally, human raters favorably assessed the overall video quality and temporal consistency of our method in a user study. Videos shown in this manuscript can be found here.
@@ 1 Introduction
Diffusion Transformers (DiTs) [35] have heralded a revolution in generative modeling and have been scaled up to generate images [5], videos [3, 9, 45, 55], and 3D assets [32] at unprecedented fidelity. Training DiTs on longer input contexts yields remarkable generalization, albeit imposing quadratic computational and memory bottlenecks. The implications are particularly apparent in the case of high-dimensional modalities like video, where training and inference with DiTs on long videos (>> 64 frames) are slow and expensive [9] (see Fig 1(a), Full Denoising). Furthermore, base models trained on short videos often show motion stagnation and repetition when extended to synthesize longer videos [2, 18]. Practitioners circumvent these constraints by generating video frames in temporal chunks using the previously generated frames as a context window for the next chunk [20, 1] proceeding in a left-to-right sequence (Fig. 1(a), Autoregressive Denoising)
@@ 3.3 Learning Objective
where $X_{t}^{i}\equiv X_{t}^{iN_{s}:(i+1)N_{s}}$ and $\epsilon_{t}^{i}\equiv\epsilon_{t}^{iN_{s}:(i+1)N_{s}}$ denote the $i^{th}$ chunk of the input and noise estimate, respectively. Eq. (8) enables the DiT to simultaneously train on different temporal chunks in parallel. Moreover, end-to-end training enables the architectures to bootstrap off each other where (1) the DiT learns to ground denoising in encoded tokens and (2) the VIN learns to refine global tokens to enable downstream denoising. In practice, neighboring chunks have useful local contexts for denoising. During training, for each chunk $i$, we also append $N_{local}$ tokens from the last $F_{local}$ frames of the previous chunk i1i-1 to the DiT.
@@ 4.6 User Study
We set up a user study (design outlined in Appendix 14) where a cohort of humans was asked to compare videos generated by VINs against other methods on (1) overall appearance and (2) temporal consistency. Fig. 12 reports the percentage results with which VIN is deemed better, comparable, or worse. In contrast to the imaging quality results of Section 4.3, humans rated VIN-generated videos to be comparable on average or even better (loss rate << 30 %). Moreover, the judged temporal consistency largely correlates with the VBench results, where VIN preference matches full generation, substantially improves upon parallel methods, and outperforms sequential generation.
@@ 5 Conclusion
Acknowledgment
This work was supported in part by an Adobe summer internship and in part by NSF under Grant No. CNS-2216746.